Jeric, a professional rooftop futbol player, kicked a πŸ’.𝟎 [π’Œπ’ˆ] ball off a rooftop that is πŸ“πŸŽ [π’Ž] off the ground, at 𝟏𝟎 [π’Ž/𝒔] oriented 30Β° degree above his horizon. His pal, Joshua, caught the ball from a window at a distant building that is πŸπŸ“ [π’Ž] off the ground. Joshua then released the ball from this height and let the ball fall to the ground. What is the speed of the ball just before it reaches Joshua? What is the speed of the ball just before it reaches the ground?


1
Expert's answer
2021-05-07T17:27:30-0400

1)


"v^2=10^2+2(9.8)(50-25)\\v=24.3\frac{m}{s}"

2)


"u^2=2(9.8)(25)\\u=22.1\frac{m}{s}"